OpenStudy (anonymous):

what is solution? write equation in standard form with steps. the info is: (5,11) m=3 and (1,-2) (-2,4)

OpenStudy (anonymous):

kat is 3/5 of the way to brianna house.larry is 7/10 of the way to briannas house. how much closer to brianna house is larry?

OpenStudy (anonymous):

larry is 1/10 closer to Brianna's house. 3/5, 7/10. 5 and 10 share a multiple. multiply 5 by 2 to equal 10 and 3 times 2 = 6/10. 7/10-6/10=1/10